what is the answer of 2+4+6....+56+58+60
Answer by stanbon(25959)

what is the answer of 2+4+6....+56+58+60
----------
Find the number of terms using a(n) = a(1) + (n-1)d:
60 = 2 + (n-1)2
58 = (n-1)2
n-1 = 29
n = 30
---------------
Find the sum of the 30 terms using S(n) = (n/2)(1st + last)
S(30) = (30/2)(2+60)
S(30) = 15*62
S(30) = 930
=================
Cheers,
Stan H.